# Higher-weight Jacobians

### v4

@@ 4.3. When is an abelian variety isomorphic to its $m$-Jacobian?
If $X$ is a $2$-maximal complex abelian variety, then by Theorem 1.2 we can assume $X=E_{1}\times\dots\times E_{n}$ is a product of $n$ elliptic curves which are CM and pairwise isogenous; by Theorem 4.1.6 we see that $X$ is $m$-maximal for all $m\leq\dim X$ and $\mathfrak{C}_{m}(X)$ is a product of $\binom{n}{m}$ pairwise isogenous CM elliptic curves. It is evident that we can only have $X\simeq\mathfrak{C}_{m}(X)$ for $m=n-1$.
Before considering for which $X=E_{1}\times E_{2}\times\dots\times E_{n}$ we have $X\simeq\mathfrak{C}_{n-1}(X)$, we will wish to check which invariants of $X$ classify it up to isomorphism. Similar to Theorem 4.2.3, where we showed that a $2$-maximal complex abelian surface $\mathcal{A}$ is uniquely classified up to isomorphism by $\mathfrak{C}(A)$ and the degree of primitivity $m$ (Remark 4.2.4) these invariants will be $\mathfrak{C}_{n}(X)$ and $n-1$ integers.
@@
We have already determined $R_{1},R_{n}$ since $r_{1}=d,r_{n}=N$. It remains to determine the $r_{i}$ for $12\leq i\leq n-$ 1.
We note that $[\mathfrak{C}_{n}(X)]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(R_{1})$ by Theorem 4.1.6. Thus by repeatedly applying Theorem 4.2.3 to $X$, we can show that
@@
where $R_{1},R_{2},R_{3}$ are the orders in $\mathcal{O}_{K}$ with conductors $d,r=f_{1}f_{2}f_{3}/(dN),N$ respectively, and $\mathfrak{C}_{3}(X)$ is an elliptic curve with CM by $R_{1}$ such that $[\mathfrak{C}_{3}(X)]=\prod_{i=1}^{3}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(R_{1})$.
@@
for $E$ an elliptic curve with CM by $R_{1}$ such that $[E]=[\mathfrak{C}_{3}(X)]^{2}\in\operatorname{Cl}(R_{1})$.
@@
It then follows that $X\simeq\mathfrak{C}(X)$ if and only if $f_{1}=f_{2}=f_{3}$ and $[E]=[\mathfrak{C}_{3}(X)]\in\operatorname{Cl}(R_{1})=\operatorname{Cl}(\mathcal{O}_{1})$; since then $[\mathfrak{C}_{3}(X)]=[E_{1}][E_{2}][E_{3}]$ and $[E]=[\mathfrak{C}_{3}(X)]^{2}$, it follows that $[\mathfrak{C}_{3}(X)]$ is the identity in $\operatorname{Cl}(\mathcal{O}_{1})$, so $\mathfrak{C}_{3}(x)\cong\mathbf{C}/\mathcal{O}_{1}$. Because in this case, $X\cong\mathbf{C}/\mathcal{O}_{1}\times\mathbf{C}/\mathcal{O}_{1}\times\mathfrak{C}_{3}(X)$, we see that $X\simeq\mathfrak{C}(X)$ is equivalent to $X\cong(\mathbf{C}/\mathcal{O}_{1})^{3}$ as we desired to show. ∎
@@
It’s computed in the proof of Proposition 4.3.3 that $X\cong\mathbf{C}/R_{3}\times\mathbf{C}/R_{2}\times\mathfrak{C}_{3}(X)$ and $\mathfrak{C}(X)\simeq\mathbf{C}/R_{2}\times\mathbf{C}/R_{1}\times E$ for $E$ an elliptic curve with CM by $R_{1}$ such that $[E]=[\mathfrak{C}_{3}(X)]^{2}\in\operatorname{Cl}(R_{1})$. A very similar computation using Theorem 4.1.5 and Theorem 4.2.3 shows that
@@
for $E^{\prime}$ an elliptic curve with CM by $R_{1}$ such that $[E^{\prime}]=[E]^{2}=[\mathfrak{C}_{3}(X)]^{4}\in\operatorname{Cl}(R_{1})$.
@@
$[E_{(k)}]=[E]^{\left(2^{k-1}\right)}=[\mathfrak{C}_{3}(X)]^{\left(2^{k}\right)}\in\operatorname{Cl}(R_{1}).$
Since $\operatorname{Cl}(R_{1})$ is finite, up to isomorphism there are only finitely many isomorphism classes $[E_{(k)}]$. Thus there are finitely many isomorphism classes of abelian threefolds of maximal Picard rank in the orbit of $X$ for the $\mathfrak{C}$ action. ∎
@@
For a $2$-maximal complex abelian variety $X$, the ($n-1$)-Jacobian $\mathfrak{C}_{n-1}(X)$ is isomorphic to $X$ if and only if $X\cong(\mathbf{C}/\mathcal{O}_{1})^{n-1}\times E$ for an elliptic curve $E$ with CM by $\mathcal{O}_{1}$ and $[E]\in\operatorname{Cl}(\mathcal{O}_{1})$ having order dividing $n-2$.
@@
where $R_{i}$ is the order in $K$ with conductor $r_{i}$, and $\mathfrak{C}_{n}(X)$ is an elliptic curve with CM by $R_{1}$ such that $[\mathfrak{C}_{n}(X)]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(R_{1})$ by Theorem 4.1.6.
@@
for $E$ an elliptic curve with CM by $R_{1}$ such that $[E]=[\mathfrak{C}_{n}(X)]^{n-1}\in\operatorname{Cl}(R_{1})$.
@@
It then follows that $X\simeq\mathfrak{C}_{n-1}(X)$ if and only if $f_{1}=\dots=f_{n}$ and $[E]=[\mathfrak{C}_{n}(X)]\in\operatorname{Cl}(R_{1})=\operatorname{Cl}(\mathcal{O}_{1})$; since $[E]=[\mathfrak{C}_{n}(X)]^{n-1}$, it follows that $[\mathfrak{C}_{n}(X)]$ has order dividing $n-2$ in $\operatorname{Cl}(\mathcal{O}_{1})$. Because in this case, $X\cong\left(\mathbf{C}/\mathcal{O}_{1}\right)^{n-1}\times\mathfrak{C}_{n}(X)$, we see that $X\simeq\mathfrak{C}_{n-1}(X)$ is equivalent to $X\cong(\mathbf{C}/\mathcal{O}_{1})^{n-1}\times E$ for an elliptic curve $E$ with CM by $\mathcal{O}_{1}$ and $[E]\in\operatorname{Cl}(\mathcal{O}_{1})$ having order dividing $n-2$, as we desired to show. ∎
@@
By Proposition 4.3.8 and Lemma 4.3.1 we see that $X=E_{1}\times\dots\times E_{n}\cong\mathfrak{C}_{n-1}(X)$ if and only if ${\rm End}(\mathfrak{C}_{n}(X))={\rm End}(E_{1})=\dots={\rm End}(E_{n})$ and $[\mathfrak{C}_{n}(X)]^{n-2}\in\operatorname{Cl}({\rm End}(\mathfrak{C}_{n}(X)))$ is the identity.
@@
It’s computed in the proof of Proposition 4.3.8 that $X\cong\mathbf{C}/R_{n}\times\dots\times\mathbf{C}/R_{2}\times\mathfrak{C}_{n}(X)$ and $\mathfrak{C}_{n-1}(X)\simeq\mathbf{C}/R_{2}\times\left(\mathfrak{C}_{n}(X)\right)^{n-1}\simeq\mathbf{C}/R_{2}\times(\mathbf{C}/R_{1})^{n-2}\times E,$ for $E$ an elliptic curve with CM by $R_{1}$ such that $[E]=[\mathfrak{C}_{n}(X)]^{n-1}\in\operatorname{Cl}(R_{1})$.
@@
for $E^{\prime}$ an elliptic curve with CM by $R_{1}$ such that $[E^{\prime}]=[E]^{n-1}=[\mathfrak{C}_{n-1}(X)]^{\left((n-1)^{2}\right)}\in\operatorname{Cl}(R_{1})$.
@@
$[E_{(k)}]=[E]^{\left((n-1)^{k-1}\right)}=[\mathfrak{C}_{3}(X)]^{\left((n-1)^{k}\right)}\in\operatorname{Cl}(R_{1}).$
Since $\operatorname{Cl}(R_{1})$ is finite, up to isomorphism there are only finitely many isomorphism classes $[E_{(k)}]$. Thus there are finitely many isomorphism classes of abelian $n$-folds of maximal Picard rank in the orbit of $X$ for the $\mathfrak{C}_{n-1}$ action. ∎
@@ 5. Kummer varieties and singular K3 surfaces
The Kummer variety can alternatively be described as the quotient of the blowup at the $2$-torsion points $\widetilde{\mathcal{A}}\to\mathcal{A}$ by a lift $\tilde{\iota}$ of the involution (cf. [huybrechts, Example 3.3]). If $\mathcal{A}$ is an abelian surface, the Kummer surface $\mathcal{K}(\mathcal{A})$ is a K3 surface.
@@
By Proposition 2.8, the blow up $\widetilde{\mathcal{A}}$ of $\mathcal{A}$ at the $2$-torsion points has an $m$-Jacobian. By Proposition 2.11 and Remark 5.2, so does $\operatorname{\mathcal{K}}(\mathcal{A})$. (The case of Brauer-Jacobians for Kummer surfaces was also shown in the proof of [si77, Theorem 6].) ∎
Corollary 5.3.
A singular complex K3 surface has a Brauer-Jacobian.
Proof.
Let $X$ be a singular K3 surface, that is, a K3 surface with Picard number $20$. By [si77], $X$ is rationally a double cover of $\operatorname{\mathcal{K}}(E\times E^{\prime})$ for two isogenous CM elliptic curves $E$ and $E^{\prime}$. Applying Corollary 2.12 gives the result. ∎
@@
In [si77], Shioda and Inose define an elliptic curve $C_{X}$ associated to a singular K3 or abelian surface $X$ which coincides with our definition of the Brauer-Jacobian $\mathfrak{C}(X)$. Furthermore, they proved (via a correspondence with equivalence classes of binary quadratic forms) that any singular complex K3 surface $Y$ (of maximal Picard rank 20) is a double cover of a Kummer surface $X=\mathcal{K}(\mathcal{A})$ associated to an abelian surface $\mathcal{A}$ of maximal Picard rank, and that $\mathfrak{C}(Y)\cong\mathfrak{C}(X)$ [si77, Theorem 4].
@@
Using the discussion of [si77], we see that any singular K3 surface $Y$ which is a double cover of the Kummer surface $\mathcal{K}(\mathcal{A})$ can be defined over the extension of the ring class field $L_{\mathcal{O}(\mathcal{A})}$ by the coordinates of the 2-torsion points of the elliptic curves $\mathbf{C}/\mathcal{O}(\mathcal{A}),\mathfrak{C}(\mathcal{A})\simeq\mathfrak{C}(Y)$, since $\mathcal{A}\cong\mathbf{C}/\mathcal{O}(\mathcal{A})\times\mathfrak{C}(Y)$ (Theorem 4.2.3) and $j(\mathfrak{C}(Y))\in L_{\mathcal{O}(\mathcal{A})}$ (Lemma 3.2.6). The field of definition of a singular K3 surface is discussed in [k3fieldofdef].
@@ Appendix A Calculating $j$-invariants
Once we have a simple description of $L_{\mathcal{O}}$, we can factor $H_{\mathcal{O}}(X)$ over $L_{\mathcal{O}}$ to get exact values of the $j$-invariants of the lattices with CM by $\mathcal{O}$. To identify which $j$-invariant corresponds to which lattice, we need only then approximate the $j$-invariants of each lattice; this can be done by truncating the $q$-series described in [wa08, Proposition 9.12].
@@
The order $\mathbf{Z}[3i]=\mathbf{Z}[\sqrt{-9}]$ in $\mathbf{Q}(i)$ has discriminant $-36$; the class group is isomorphic to $\mathbf{Z}/2\mathbf{Z}$, with the two classes being represented by the quadratic forms $x^{2}+9y^{2},2x^{2}+2xy+5y^{2}$ or the lattices $\langle 1,3i\rangle,\langle 2,-1+3i\rangle\sim\langle 3,1+i\rangle$, with $j$-invariants
@@
The order $\mathbf{Z}[6i]=\mathbf{Z}[\sqrt{-36}]$ in $\mathbf{Q}(i)$ has discriminant $-144$; the class group is isomorphic to $\mathbf{Z}/4\mathbf{Z}$, with the four classes being represented by the quadratic forms (and corresponding lattices)
@@
[…]
@@
The order $\mathbf{Z}[3\sqrt{-3}]=\mathbf{Z}[\sqrt{-27}]$ in $\mathbf{Q}(\sqrt{-3})$ has discriminant $-27$; the class group is isomorphic to $\mathbf{Z}/3\mathbf{Z}$, with the three classes being represented by the quadratic forms (and corresponding lattices)
@@
[…]
@@
The order $\mathbf{Z}[4\sqrt{-3}]=\mathbf{Z}[\sqrt{-48}]$ in $\mathbf{Q}(\sqrt{-3})$ has discriminant $-48$; the class group is isomorphic to $\mathbf{Z}/2\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$, with the four classes being represented by the quadratic forms (and corresponding lattices)
@@
[…]

### v5

@@ 4.3. When is an abelian variety isomorphic to its $m$-Jacobian?
If $X$ is a $2$-maximal complex abelian variety, then by Theorem 1.2 we can assume $X=E_{1}\times\dots\times E_{n}$ is a product of $n$ elliptic curves which are CM and pairwise isogenous; by Theorem 4.1.6 we see that $X$ is $m$-maximal for all $m\leq\dim X$ and $\mathfrak{C}_{m}(X)$ is a product of $\binom{n}{m}$ pairwise isogenous CM elliptic curves. It is evident that we can only have $X\simeq\mathfrak{C}_{m}(X)$ for m=n1m=n-1.
Before considering for which $X=E_{1}\times E_{2}\times\dots\times E_{n}$ we have $X\simeq\mathfrak{C}_{n-1}(X)$, we will wish to check which invariants of $X$ classify it up to isomorphism. Similar to Theorem 4.2.3, where we showed that a $2$-maximal complex abelian surface $\mathcal{A}$ is uniquely classified up to isomorphism by $\mathfrak{C}(A)$ and the degree of primitivity $m$ (Remark 4.2.4) these invariants will be $\mathfrak{C}_{n}(X)$ and n1n-1 integers.
@@
We have already determined $R_{1},R_{n}$ since $r_{1}=d,r_{n}=N$. It remains to determine the $r_{i}$ for $2\leq i\leq n-1$.
We note that $\mathfrak{C}_{n}(X)]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(R_{1})$ by Theorem 4.1.6. Thus by repeatedly applying Theorem 4.2.3 to $X$, we can show that
@@
where $R_{1},R_{2},R_{3}$ are the orders in $\mathcal{O}_{K}$ with conductors $d,r=f_{1}f_{2}f_{3}/(dN),N$ respectively, and $\mathfrak{C}_{3}(X)$ is an elliptic curve with CM by $R_{1}$ such that $\mathfrak{C}_{3}(X)]=\prod_{i=1}^{3}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(R_{1})$.
@@
for $E$ an elliptic curve with CM by $R_{1}$ such that $E]=[\mathfrak{C}_{3}(X)]^{2}\in\operatorname{Cl}(R_{1})$.
@@
It then follows that $X\simeq\mathfrak{C}(X)$ if and only if $f_{1}=f_{2}=f_{3}$ and $E]=[\mathfrak{C}_{3}(X)]\in\operatorname{Cl}(R_{1})=\operatorname{Cl}(\mathcal{O}_{1})$; since then $\mathfrak{C}_{3}(X)]=[E_{1}][E_{2}][E_{3}]$ and $E]=[\mathfrak{C}_{3}(X)]^{2}$, it follows that $\mathfrak{C}_{3}(X)]$ is the identity in $\operatorname{Cl}(\mathcal{O}_{1})$, so $\mathfrak{C}_{3}(X)\cong\mathbf{C}/\mathcal{O}_{1}$. Because in this case, $X\cong\mathbf{C}/\mathcal{O}_{1}\times\mathbf{C}/\mathcal{O}_{1}\times\mathfrak{C}_{3}(X)$, we see that $X\simeq\mathfrak{C}(X)$ is equivalent to $X\cong(\mathbf{C}/\mathcal{O}_{1})^{3}$ as we desired to show. ∎
@@
It’s computed in the proof of Proposition 4.3.3 that $X\cong\mathbf{C}/R_{3}\times\mathbf{C}/R_{2}\times\mathfrak{C}_{3}(X)$ and $\mathfrak{C}(X)\simeq\mathbf{C}/R_{2}\times\mathbf{C}/R_{1}\times E$ for $E$ an elliptic curve with CM by $R_{1}$ such that $E]=[\mathfrak{C}_{3}(X)]^{2}\in\operatorname{Cl}(R_{1})$. A very similar computation using Theorem 4.1.5 and Theorem 4.2.3 shows that
@@
for $E^{\prime}$ an elliptic curve with CM by $R_{1}$ such that $E^{\prime}]=[E]^{2}=[\mathfrak{C}_{3}(X)]^{4}\in\operatorname{Cl}(R_{1})$.
@@
$E_{(k)}]=[E]^{\left(2^{k-1}\right)}=[\mathfrak{C}_{3}(X)]^{\left(2^{k}\right)}\in\operatorname{Cl}(R_{1}).$
Since $\operatorname{Cl}(R_{1})$ is finite, up to isomorphism there are only finitely many isomorphism classes $E_{(k)}]$. Thus there are finitely many isomorphism classes of abelian threefolds of maximal Picard rank in the orbit of $X$ for the $\mathfrak{C}$ action. ∎
@@
For a $2$-maximal complex abelian variety $X$, the (n1n-1)-Jacobian $\mathfrak{C}_{n-1}(X)$ is isomorphic to $X$ if and only if $X\cong(\mathbf{C}/\mathcal{O}_{1})^{n-1}\times E$ for an elliptic curve $E$ with CM by $\mathcal{O}_{1}$ and $E]\in\operatorname{Cl}(\mathcal{O}_{1})$ having order dividing n2n-2.
@@
where $R_{i}$ is the order in $K$ with conductor $r_{i}$, and $\mathfrak{C}_{n}(X)$ is an elliptic curve with CM by $R_{1}$ such that $\mathfrak{C}_{n}(X)]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(R_{1})$ by Theorem 4.1.6.
@@
for $E$ an elliptic curve with CM by $R_{1}$ such that $E]=[\mathfrak{C}_{n}(X)]^{n-1}\in\operatorname{Cl}(R_{1})$.
@@
It then follows that $X\simeq\mathfrak{C}_{n-1}(X)$ if and only if $f_{1}=\dots=f_{n}$ and $E]=[\mathfrak{C}_{n}(X)]\in\operatorname{Cl}(R_{1})=\operatorname{Cl}(\mathcal{O}_{1})$; since $E]=[\mathfrak{C}_{n}(X)]^{n-1}$, it follows that $\mathfrak{C}_{n}(X)]$ has order dividing n2n-2 in $\operatorname{Cl}(\mathcal{O}_{1})$. Because in this case, $X\cong\left(\mathbf{C}/\mathcal{O}_{1}\right)^{n-1}\times\mathfrak{C}_{n}(X)$, we see that $X\simeq\mathfrak{C}_{n-1}(X)$ is equivalent to $X\cong(\mathbf{C}/\mathcal{O}_{1})^{n-1}\times E$ for an elliptic curve $E$ with CM by $\mathcal{O}_{1}$ and $E]\in\operatorname{Cl}(\mathcal{O}_{1})$ having order dividing n2n-2, as we desired to show. ∎
@@
By Proposition 4.3.8 and Lemma 4.3.1 we see that $X=E_{1}\times\dots\times E_{n}\cong\mathfrak{C}_{n-1}(X)$ if and only if ${\rm End}(\mathfrak{C}_{n}(X))={\rm End}(E_{1})=\dots={\rm End}(E_{n})$ and $\mathfrak{C}_{n}(X)]^{n-2}\in\operatorname{Cl}({\rm End}(\mathfrak{C}_{n}(X)))$ is the identity.
@@
It’s computed in the proof of Proposition 4.3.8 that $X\cong\mathbf{C}/R_{n}\times\dots\times\mathbf{C}/R_{2}\times\mathfrak{C}_{n}(X)$ and $\mathfrak{C}_{n-1}(X)\simeq\mathbf{C}/R_{2}\times\left(\mathfrak{C}_{n}(X)\right)^{n-1}\simeq\mathbf{C}/R_{2}\times(\mathbf{C}/R_{1})^{n-2}\times E,$ for $E$ an elliptic curve with CM by $R_{1}$ such that $E]=[\mathfrak{C}_{n}(X)]^{n-1}\in\operatorname{Cl}(R_{1})$.
@@
for $E^{\prime}$ an elliptic curve with CM by $R_{1}$ such that $E^{\prime}]=[E]^{n-1}=[\mathfrak{C}_{n-1}(X)]^{\left((n-1)^{2}\right)}\in\operatorname{Cl}(R_{1})$.
@@
$E_{(k)}]=[E]^{\left((n-1)^{k-1}\right)}=[\mathfrak{C}_{3}(X)]^{\left((n-1)^{k}\right)}\in\operatorname{Cl}(R_{1}).$
Since $\operatorname{Cl}(R_{1})$ is finite, up to isomorphism there are only finitely many isomorphism classes $E_{(k)}]$. Thus there are finitely many isomorphism classes of abelian $n$-folds of maximal Picard rank in the orbit of $X$ for the $\mathfrak{C}_{n-1}$ action. ∎
@@ 5. Kummer varieties and singular K3 surfaces
The Kummer variety can alternatively be described as the quotient of the blowup at the $2$-torsion points $\widetilde{\mathcal{A}}\to\mathcal{A}$ by a lift $\tilde{\iota}$ of the involution (cf. [Huy16, Example 3.3]). If $\mathcal{A}$ is an abelian surface, the Kummer surface $\mathcal{K}(\mathcal{A})$ is a K3 surface.
@@
[…]
The involution $\iota$ has a trivial action on the even degree cohomology of $\mathcal{A}$. Thus $\mathfrak{C}_{2m}(\operatorname{\mathcal{K}}(\mathcal{A}))\cong\mathfrak{C}_{2m}(\mathcal{A})$ as we desired to show. (The case of Brauer-Jacobians for Kummer surfaces was also shown in the proof of [SI77, Theorem 6]. Moreover, for any $q$ odd we have $\mathfrak{C}_{q}(\operatorname{\mathcal{K}}(\mathcal{A}))=0$ by Proposition 2.12.) ∎
@@
In [SI77], Shioda and Inose define an elliptic curve $C_{X}$ associated to a singular K3 or abelian surface $X$ which coincides with our definition of the Brauer-Jacobian $\mathfrak{C}(X)$. Furthermore, they proved (via a correspondence with equivalence classes of binary quadratic forms) that any singular complex K3 surface $Y$ (of maximal Picard rank 20) is a double cover of a Kummer surface $X=\mathcal{K}(\mathcal{A})$ associated to an abelian surface $\mathcal{A}$ of maximal Picard rank, and that $\mathfrak{C}(Y)\cong\mathfrak{C}(X)$ [SI77, Theorem 4]. Therefore to compute higher-weight Jacobians for complex K3 surfaces in general, it is enough to do so for Kummer surfaces.
@@
Using the Shioda-Inose construction of [SI77], Schuett proves that a singular K3 surface $Y$ which is a double cover of a Kummer surface $\mathcal{K}(E_{1}\times E_{2})$ can be defined over the field $\mathbf{Q}(j(E_{1}),j(E_{2}))$ [Sch06, Proposition 10, Remark 11]. The field of definition of a singular K3 surface associated to an intersection form (or elliptic curve $C_{X}$) with degree of primitivity $m=1,2,3$ and certain discriminant values is discussed in [Sch06, Lemma 20]. In Proposition 1.2 we consider the field of definition of a primitive $(m=1)$ abelian surface (and hence its Kummer surface) compared to the field of definition of its Brauer-Jacobian for any discriminant.
@@ Appendix A Calculating $j$-invariants
Once we have a simple description of $L_{\mathcal{O}}$, we can factor $H_{\mathcal{O}}(X)$ over $L_{\mathcal{O}}$ to get exact values of the $j$-invariants of the lattices with CM by $\mathcal{O}$. To identify which $j$-invariant corresponds to which lattice, we need only then approximate the $j$-invariants of each lattice; this can be done by truncating the $q$-series described in [Was08, Proposition 9.12].
@@
The order $\mathbf{Z}[3i]=\mathbf{Z}[\sqrt{-9}]$ in $\mathbf{Q}(i)$ has discriminant 36-36; the class group is isomorphic to $\mathbf{Z}/2\mathbf{Z}$, with the two classes being represented by the quadratic forms $x^{2}+9y^{2},2x^{2}+2xy+5y^{2}$ or the lattices $\langle 1,3i\rangle,\langle 2,-1+3i\rangle\sim\langle 3,1+i\rangle$, with $j$-invariants
@@
The order $\mathbf{Z}[6i]=\mathbf{Z}[\sqrt{-36}]$ in $\mathbf{Q}(i)$ has discriminant 144-144; the class group is isomorphic to $\mathbf{Z}/4\mathbf{Z}$, with the four classes being represented by the quadratic forms (and corresponding lattices)
@@
$j(\langle 1,6i\rangle)$ $=5894625992142600+3403263903336192\sqrt{3}$
$+3167093925247392\sqrt[4]{12}+914261265145368\left(\sqrt[4]{12}\right)^{3},$
$j(\langle 3,1-2i\rangle)$ $=5894625992142600-3403263903336192\sqrt{3}$
$-i\left(3167093925247392\sqrt[4]{12}-914261265145368\left(\sqrt[4]{12}\right)^{3}\right),$
$j(\langle 3,1+2i\rangle)$ $=5894625992142600-3403263903336192\sqrt{3}$
$+i\left(3167093925247392\sqrt[4]{12}-914261265145368\left(\sqrt[4]{12}\right)^{3}\right),$
$j(\langle 3,2i\rangle)$ $=5894625992142600+3403263903336192\sqrt{3}$
$-3167093925247392\sqrt[4]{12}-914261265145368\left(\sqrt[4]{12}\right)^{3}.$
@@
The order $\mathbf{Z}[3\sqrt{-3}]=\mathbf{Z}[\sqrt{-27}]$ in $\mathbf{Q}(\sqrt{-3})$ has discriminant 27-27; the class group is isomorphic to $\mathbf{Z}/3\mathbf{Z}$, with the three classes being represented by the quadratic forms (and corresponding lattices)
@@
$j(\langle 1,3\sqrt{-3}\rangle)$ $=31710790944000\left(\sqrt[3]{2}\right)^{2}+39953093016000\sqrt[3]{2}+50337742902000,$
$j(\langle 3,2+\sqrt{-3}\rangle)$ $=31710790944000\left(\zeta_{3}\sqrt[3]{2}\right)^{2}+39953093016000\zeta_{3}\sqrt[3]{2}+50337742902000,$
$j(\langle 3,1+\sqrt{-3}\rangle)$ $=31710790944000\left(\zeta_{3}^{2}\sqrt[3]{2}\right)^{2}+39953093016000\zeta_{3}^{2}\sqrt[3]{2}+50337742902000.$
@@
The order $\mathbf{Z}[4\sqrt{-3}]=\mathbf{Z}[\sqrt{-48}]$ in $\mathbf{Q}(\sqrt{-3})$ has discriminant 48-48; the class group is isomorphic to $\mathbf{Z}/2\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$, with the four classes being represented by the quadratic forms (and corresponding lattices)
@@
$j(\langle 1,4\sqrt{-3}\rangle)$ $=820762881440077125\sqrt{6}+1160733998424384000\sqrt{3}$
$+1421603011620136125\sqrt{2}+2010450259344609000,$
$j(\langle 4,2+\sqrt{-3}\rangle)$ $=-820762881440077125\sqrt{6}-1160733998424384000\sqrt{3}$
$+1421603011620136125\sqrt{2}+2010450259344609000,$
$j(\langle 2,1+2\sqrt{-3}\rangle)$ $=-820762881440077125\sqrt{6}+1160733998424384000\sqrt{3},=$
$-1421603011620136125\sqrt{2}+2010450259344609000,$
$j(\langle 4,\sqrt{-3}\rangle)$ $=820762881440077125\sqrt{6}-1160733998424384000\sqrt{3}$
$-1421603011620136125\sqrt{2}+2010450259344609000.$